OpenStudy (anonymous):

how do you solve r/4-r/14=1/4

OpenStudy (bahrom7893):

multiply through by 14

OpenStudy (bahrom7893):

wait nevermind not 14

OpenStudy (bahrom7893):

multiply through by 56 to get: 14r-4r=14

OpenStudy (bahrom7893):

10r=14 r=14/10=7/5